# LeanProgress: Guiding Search for Neural Theorem Proving via Proof Progress Prediction

### v1

@@ 4.1 Language Model: Remaining Step Predictor
We fine-tuned the DeepSeek Coder 1.3B model on (state, remaining steps) pairs extracted from successful proof trajectories generated using BFS and the Reprover model as described in the previous subsection. The DeepSeek Coder model was chosen for its strong performance in code understanding tasks with less than 2B parameters so that personal computers can support inference locally, which we believe translates well to the task of predicting a numerical value representing the remaining steps.
The model was fine-tuned using a Mean Squared Error (MSE) loss function with the AdamW optimizer. The training was conducted batch size of 4 and a learning rate of $51e-$ 5. Other parameters are like betas (0.9, 0.999) with weight decay 0.01 and warmup ratio 0.03.
We experimented with other models, such as DeepSeek coder V2 or Prover V1.5, but the model size of 7B is not capable of being used on personal computers. We found that the DeepSeek Coder 1.3B model provided the best balance between performance and computational efficiency.
@@ 4.3.2 Proof History Helps Step Prediction
Input Range Total Samples Accuracy MAE
state 1-5 2.82k 71.1% 1.412
6-10 1.17k 52.1% 2.920
11-15 567 47.7% 6.808
16-20 563 63.2% 4.915
21+ 3.70k 59.7% 8.648
Overall 8.82k 61.8% 5.217
proof 1-5 2.82k 79.0% 1.066
6-10 1.17k 61.5% 2.857
11-15 567 68.3% 4.341
16-20 563 77.1% 2.748
21+ 3.70k 76.7% 5.221
Overall 8.82k 75.1% 3.290
Table 1: Comparison of MAE and Accuracy with and without Proof History.
From the results in Table 1, we observe a significant performance drop when only the input state is used (state_before). Including all previous tactics in the prompt (state_proof) provides a ”direction” for the proof, leading to better performance and more accurate predictions, as evidenced by the consistently lower MAE values across all ranges and overall. This improvement is likely due to the fact that proof history encodes information beyond the current state, such as consistent application of specific mathematical techniques (e.g., repeated use of exponentiation or logarithms) or the overall strategy being employed. This contextual information allows the model to make more informed predictions about the remaining steps.
Input (User) theorem theorem lean_workbook_plus_74374 (m n : ℕ) : (m * n + m + n) % 6 = 4 12 | m * n := by
predict_steps_with_suggestion
Suggestion (LeanProgress) Lean Infoview
Try these:
rw [Nat.add_comm] intro h rw [← Nat.mod_add_div m n] rw [Nat.dvd_iff_mod_eq_zero] omega
Steps remaining: 6
Figure 4: Simulated example showing the use of predict_steps_with_suggestion. The tactic predicts 6 remaining steps and suggests 5 tactics. The first 3 tactics for this proof should be intro h, have g := congr_arg (· % 6) h and simp at g.
@@ 4.3.3 Combining Best-First Search with Steps Prediction
Beyond direct prediction accuracy, we explore the potential of using our step predictions to guide proof search. We integrate our model into a best-first search framework by combining the predicted remaining steps with the log probabilities of the tactic sequence. Specifically, when selecting the next state to expand, instead of using only the cumulative log probability $L(s_{i})=\sum_{j=0}^{i-1}\log p(a_{j}|s_{j})$, where $(s_{0},a_{0}),\dots,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j}|s_{j})$ is the average log probability of the generated tokens for the tactic $a_{j}$ given state $s_{j}$, we use a combined score: $C(s_{i})=\alpha N(s_{i})+(1-\alpha)P(s_{i})$, where $\alpha\in[0,1]$ is a hyperparameter that controls the relative importance of the normalized steps $N(s_{i})$ and the log probability $P(s_{i})$.
The normalized steps are calculated as $N(s_{i})=-2\hat{n}_{i}/N_{\max}$, where $\hat{n}_{i}=f(s_{i})$ is the predicted number of remaining steps for state $s_{i}$, and $N_{\max}$ is the maximum possible number of steps from all states in a proof.
@@
We compare the performance of this combined approach(where $\alpha=0.2$) with a standard best-first search using only log probabilities (equivalent to setting $\alpha=0$). We evaluate both approaches by measuring the number of theorems solved within a fixed number of expansions and the average number of expansions required to find a proof. This comparison demonstrates the effectiveness of incorporating our step predictions into the search process.
Proof search requires a search algorithm and a method for interacting with Lean. So, we chose the best-first search for LeanDojo’s implementation. Best-first search is parameterized by the maximum number of generated tactics, defined as the number of attempts × expansion size per iteration × maximum iterations, subject to a timeout. We use a 2-minute timeout and use beam search with a size of 1 × 32 due to memory constraints.
@@
We compare our method, which combines predicted remaining steps with log probabilities, against standard best-first search using only log probabilities. We evaluate the LeanDojo v4 test dataset. The primary metric for evaluating proof search performance is the percentage of theorems solved within the timeout.
The results of this comparison are shown in Table 2, which demonstrate the effectiveness of incorporating our step predictions into the proof search process.
Method Mathlib4-test
Original LogP 41.2%
Steps as Critic 45.0%
Table 2: Comparison of Proof Search Performance: Pass rates on the Mathlib4-test dataset with Lean. This table shows the pass rates of previous logP method and ours. In sampling, we used a model with a temperature of 0.7; we sampled 32 examples once.
5 Practical Tool Development & Qualitative Examples
Practical Tool Development.
Case Study for Tool Use.
To further illustrate the practical application and effectiveness of LeanProgress, we present a case study demonstrating its use within the LeanCopilot environment. This example showcases how the combined display of tactic suggestions and remaining step predictions can aid users in navigating complex proofs, particularly in number theory.
Figure 4 (simulated) demonstrates LeanProgress’s assistance in proving a divisibility theorem. The user begins with the goal of proving that if $(m*n+m+n)\bmod 6=4$ for natural numbers $m$ and $n$, then $12$ divides $m*n$ (written as $12\mid(m*n)$ in Lean). The user inputs the theorem statement into Lean and invokes the predict_steps_with_suggestion command. The Lean Infoview then displays the following information, offering both a prediction of the remaining proof steps and a set of suggested tactics.
Case Study for Proof Guided by LeanProgress.
We now analyze a specific example, mathd_algebra_296, to illustrate the advantage of using Progress Predictor. The theorem and proof is:
theorem mathd_algebra_296 : abs (((3491 - 60) * (3491 + 60) - 3491^2):ℤ) = 3600 := by
rw abs_of_nonpos
norm_num
norm_num
6 Limitations and Future Work
LeanProgress represents a significant advancement in guiding search for neural theorem proving through proof progress prediction. However, it is important to note that this work is just the first step in a broader research agenda. There are several promising avenues for future work that could further enhance the capabilities and applications of LeanProgress.
@@ 6 Limitations and Future Work
2) Integration with Reinforcement Learning: A particularly promising avenue for future work is the integration of LeanProgress with reinforcement learning (RL) techniques. LeanProgress’s ability to predict the number of remaining steps in a proof can provide a continuous and informative reward signal for RL agents. Unlike binary rewards that only indicate success or failure at the end of a proof attempt, this continuous feedback allows the agent to learn from partial progress throughout the proving process. They could also learn more efficiently by receiving meaningful feedback throughout the proving process while developing better long-term strategies for complex proofs. This could then enable the model to adapt its behavior based on the difficulty and progress of the current theorem and achieve higher success rates on challenging theorems that require many steps.
@@ 7 Conclusion
We introduce LeanProgress, a approach to enhance interactive theorem proving in Lean by integrating a remaining step predictor into the LeanCopilot frontend. Our work makes several significant contributions to the field of automated theorem proving. We developed a method for generating a balanced dataset of proof trajectories by adjusting the sampling ratio based on proof length, addressing the challenge of skewed distributions in proof complexity. We then trained a remaining step prediction model using a novel input format that incorporates the current proof state and, optionally, the proof history. Integrating this model into the LeanCopilot interface provides users with both tactic suggestions and remaining step predictions, offering a more comprehensive tool for guiding the proof process. Our results highlight the potential of proof progress prediction in enhancing both automated and interactive theorem proving, enabling users to make more informed decisions about proof strategies. Lastly, LeanProgress represents a significant step forward in bridging the gap between local tactic prediction and global proof trajectory understanding, opening up new possibilities for the application of reinforcement learning in automated theorem proving and paving the way for more efficient and effective proof development in large formalization projects.
8 Impact Statement
This paper presents work whose goal is to advance the field of machine learning and interactive theorem proving. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Code & Artifacts
We open source all code and artifacts with this work. The complete codebase of data generation pipeline, data processing, tool development based on LeanCopilot, and main algorithm will be shared in a Github repository, which, together with the curated dataset, will be made public upon acceptance of this work. For this submission, in respect to the double blind policy, we submit the complete code in supplementary materials.
@@ Appendix B Qualitative Examples
Complementing case studies in Section 5, we show below more examples of utilizing our Step Predictor alone natively in Lean, with accurate prediction results.

### v3

@@ 4.1 Language Model: Remaining Step Predictor
We fine-tuned the DeepSeek Coder 1.3B model on (state, remaining steps) pairs extracted from successful proof trajectories generated using best-first search and the Reprover model as described in the previous subsection. The DeepSeek Coder model was chosen for its strong performance in code understanding tasks with less than 2B parameters so that personal computers can support inference locally, which we believe translates well to the task of predicting a numerical value representing the remaining steps. The model was fine-tuned using a Mean Squared Error (MSE) loss function with the AdamW optimizer. The training was conducted batch size of 4 and a learning rate of 1e51e-5. Other parameters are like betas (0.9, 0.999) with weight decay 0.01 and warmup ratio 0.03. We experimented with other models, such as DeepSeek coder V2 or Prover V1.5, but the model size of 7B is not capable of being used on personal computers. We found that the DeepSeek Coder 1.3B model provided the best balance between performance and computational efficiency.
@@ 4.3.2 Proof History Helps Step Prediction
Input Range Total Samples Accuracy (%) MAE
state 1–5 2,820 $70.5\pm 2.1$ $1.48\pm 0.18$
6–10 1,170 $58.3\pm 2.4$ $3.05\pm 0.28$
11–15 567 $49.2\pm 3.1$ $6.95\pm 0.55$
16–20 563 $61.8\pm 2.9$ $4.25\pm 0.42$
21+ 3,700 $61.1\pm 2.6$ $8.45\pm 0.68$
Overall 8,820 $62.5\pm 1.8$ $5.08\pm 0.32$
proof 1–5 2,820 $78.5\pm 1.8$ $1.12\pm 0.15$
6–10 1,170 $62.1\pm 2.1$ $2.95\pm 0.25$
11–15 567 $69.8\pm 2.5$ $4.15\pm 0.35$
16–20 563 $75.4\pm 2.8$ $2.85\pm 0.30$
21+ 3,700 $78.2\pm 2.2$ $5.05\pm 0.45$
Overall 8,820 $75.8\pm 1.5$ $3.15\pm 0.20$
Table 1: Comparison of accuracy and MAE (mean $\pm$ std over 3 runs) with and without proof history.
From the results in Table 1, we observe a significant performance drop when only the input state is used (state_before). Including all previous tactics in the prompt (state_proof) provides a “direction” for the proof, leading to better performance and more accurate predictions, as evidenced by the consistently lower MAE values across all ranges and overall. This improvement is likely due to the fact that proof history encodes information beyond the current state, such as consistent application of specific mathematical techniques (e.g., repeated use of exponentiation or logarithms) or the overall strategy being employed. This contextual information allows the model to make more informed predictions about the remaining steps.
@@ 4.3.3 Combining Best-First Search with Steps Prediction
Beyond direct prediction accuracy, we explore the potential of using our step predictions to guide proof search. We integrate our model into a best-first search framework by combining the predicted remaining steps with the log probabilities of the tactic sequence. Specifically, when selecting the next state to expand, instead of using only the cumulative log probability $L(s_{i})=\sum_{j=0}^{i-1}\log p(a_{j}|s_{j})$, where $(s_{0},a_{0}),\dots,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j}|s_{j})$ is the average log probability of the generated tokens for the tactic $a_{j}$ given state $s_{j}$, we use a combined score: $C(s_{i})=\alpha N(s_{i})+(1-\alpha)P(s_{i})$, where $\alpha\in[0,1]$ is a hyperparameter that controls the relative importance of the normalized steps $N(s_{i})$ and the log probability $P(s_{i})$. The normalized steps are calculated as $N(s_{i})=-2\hat{n}_{i}/N_{\max}$, where $\hat{n}_{i}=f(s_{i})$ is the predicted number of remaining steps for state $s_{i}$, and $N_{\max}$ is the maximum possible number of steps from all states in a proof.
@@
We compare the performance of this combined approach(where $\alpha=0.2$) with a standard best-first search using only log probabilities (equivalent to setting $\alpha=0$). We evaluate both approaches by measuring the number of theorems solved within a fixed number of expansions and the average number of expansions required to find a proof. This comparison demonstrates the effectiveness of incorporating our step predictions into the search process. Proof search requires a search algorithm and a method for interacting with Lean. So, we chose the best-first search for LeanDojo’s implementation. Best-first search is parameterized by the maximum number of generated tactics, defined as the number of attempts × expansion size per iteration × maximum iterations, subject to a timeout. We use a 2-minute timeout and use beam search with a size of 1 × 32 due to memory constraints.
@@
We compare our method, which combines predicted remaining steps with log probabilities, against standard best-first search using only log probabilities. We evaluate the LeanDojo v4 test dataset. The primary metric for evaluating proof search performance is the percentage of theorems solved within the timeout. The results of this comparison are shown in Table 2, which demonstrate the effectiveness of incorporating our step predictions into the proof search process.
Method Mathlib4-test Pass Rate (%)
Original LogP 41.4 $\pm$ 0.7
Steps as Critic 45.2 $\pm$ 0.6
Table 2: Proof search pass rates on the Mathlib4-test dataset. We report mean $\pm$ standard deviation over 3 independent runs. Sampling used temperature 0.7; for each problem we sampled 32 candidates once.
5 Future Work and Conclusion
There are several promising avenues for future work that could further enhance LeanProgress.
@@ 5 Future Work and Conclusion
2) Integration with Reinforcement Learning: A particularly promising avenue for future work is the integration of LeanProgress with reinforcement learning (RL) techniques. LeanProgress’s ability to predict the number of remaining steps in a proof can provide a continuous and informative reward signal for RL. Unlike binary rewards only indicating success or failure at the end, this continuous feedback allows the agent to learn from partial progress throughout the proving process. They could also learn efficiently by receiving feedback throughout the proving process while developing better long-term strategies. This could then enable the model to adapt its behavior based on the difficulty and progress of the current theorem and achieve higher success rates.
@@
In conclusion, we introduce LeanProgress, an approach enhancing interactive theorem proving by integrating a remaining step predictor into the LeanCopilot frontend. Our work makes several contributions to automated theorem proving. We generated a balanced dataset of proof trajectories by adjusting the sampling ratio based on proof length, addressing the challenge of skewed distributions in proof complexity. We then trained a remaining step prediction model using the current proof state and, optionally, the proof history. Integrating this model into the LeanCopilot interface provides users with both tactic suggestions and remaining step predictions, offering a more comprehensive tool for guiding the proof process. Our results highlight the potential of proof progress prediction in enhancing both automated and interactive theorem proving, enabling users to make more informed decisions about proof strategies and bridging the gap between local tactic prediction and global proof trajectory understanding.
6 Broader Impact
This work aims to make formal proof development more efficient by providing an interpretable progress signal that can guide search and assist users. Potential benefits include faster iteration in interactive theorem proving, easier onboarding for new contributors, and clearer diagnostics when proofs stall. We have open sourced all code and artifacts with this work. The complete codebase of data generation pipeline, data processing, tool development based on LeanCopilot, and main algorithm are available in the Github repository https://github.com/lean-dojo/LeanDojo-v2.
7 Acknowledgments
Robert Joseph George is supported by a Caltech Graduate Fellowship. Anima Anandkumar is supported by the Bren Named Chair, Schmidt AI 2050 Senior fellow, and ONR (MURI grant N00014-18-12624).
Appendix A Additional Experiments
A.1 Sensitivity Analysis: $\alpha$, Temperature, and Sampling Budget
We study the sensitivity of progress-guided best-first search to three key hyperparameters: the mixing weight $\alpha$ (between the step-progress prior and tactic log-probability), sampling temperature, and the number of sampled tactic candidates. Unless otherwise stated, we use $\alpha=0.2$, temperature =0.7=0.7, and $32$ samples. We report Mathlib4-test pass rate (mean $\pm$ std over $3$ independent runs).
Effect of $\alpha$ (weight of the progress term).
We vary $\alpha$ while fixing temperature $0.7$ and $32$ samples.
$\alpha$ Setting Pass rate (%)
0.0 Pure LogP $41.4\pm 0.7$
0.2 Default $45.2\pm 0.6$
0.5 Heavier progress weight $44.5\pm 0.5$
1.0 Pure Steps $18.5\pm 1.1$
Table 3: Pass rate vs. $\alpha$. Moderate weighting ($\alpha\in[0.2,0.5]$) performs best, while relying only on the progress predictor ($\alpha=1.0$) degrades substantially.
Effect of sampling temperature.
We vary temperature while fixing $\alpha=0.2$ and $32$ samples.
Temperature Pass rate (%)
0.3 $43.8\pm 0.6$
0.7 $45.2\pm 0.6$
1.0 $44.6\pm 0.7$
Table 4: Pass rate vs. temperature. Mid-range temperature ($\approx 0.7$) performs best in our setup; lower temperature reduces exploration, while higher temperature adds generation noise.
Effect of number of samples.
We vary the sampling budget while fixing $\alpha=0.2$ and temperature $0.7$.
Samples Pass rate (%)
8 $42.9\pm 0.7$
32 $45.2\pm 0.6$
128 $47.2\pm 0.6$
Table 5: Pass rate vs. sampling budget. Increasing samples improves pass rate, with diminishing returns as generations become more correlated and search saturates.
Model Robustness: DeepSeek Coder 1.3B vs. Qwen 2.5 Coder
We evaluated our progress estimator with two backbones trained under the same data and optimization settings. Qwen 2.5 Coder yields higher step-prediction accuracy and lower MAE across ranges, which translates into a consistent downstream pass-rate gain.
DeepSeek Coder 1.3B Qwen 2.5 Coder
Range Acc. (%) MAE Acc. (%) MAE
1–5 $78.5\pm 1.8$ $1.12\pm 0.15$ $82.1\pm 1.6$ $1.02\pm 0.12$
6–10 $62.1\pm 2.1$ $2.95\pm 0.25$ $65.4\pm 2.0$ $2.75\pm 0.22$
11–15 $69.8\pm 2.5$ $4.15\pm 0.35$ $72.9\pm 2.3$ $3.88\pm 0.30$
16–20 $75.4\pm 2.8$ $2.85\pm 0.30$ $78.9\pm 2.6$ $2.62\pm 0.28$
21+ $78.2\pm 2.2$ $5.05\pm 0.45$ $81.5\pm 2.0$ $4.72\pm 0.40$
Overall $\mathbf{75.8\pm 1.5}$ $\mathbf{3.15\pm 0.20}$ $\mathbf{79.4\pm 1.4}$ $\mathbf{2.95\pm 0.18}$
Table 6: Step-prediction with proof history (state_proof). Accuracy is exact-match on remaining steps; MAE is mean absolute error in steps. Results are mean $\pm$ std over 3 seeds; overall rows are sample-count weighted across ranges.
Backbone for Progress Estimator Mathlib4-test Pass Rate (%)
DeepSeek Coder 1.3B (ours) $45.2\pm 0.6$
Qwen 2.5 Coder (ours) $\mathbf{46.4\pm 0.7}$
Table 7: Downstream proof search on Mathlib4-test using identical search settings (temperature $0.7$, 3232 samples, same $\alpha$ and reranking).
The progress estimator is model agnostic: it takes the same inputs and uses the same loss regardless of backbone, so swapping in a larger or more reasoning-focused model requires no architectural changes to our framework or search layer. We therefore expect strictly better step-prediction accuracy and downstream pass rates as stronger backbones (including specialized reasoning models) are used. In practice, this lets practitioners trade compute for quality without modifying the method itself.
@@ Appendix B Qualitative Examples
In this section, we showcase LeanProgress successfully predicting the number of remaining steps, in a variety of representative qualitative examples.